OpenStudy (anonymous):

Solve: 4(x + 1) = −4

OpenStudy (mayankdevnani):

4x+4=-4 4x=-8 x=-2

OpenStudy (jiteshmeghwal9):

distribut 4 btw x+1 4x+4=-4 4x=-4-4 4x=-8 x=-8/4=-2

OpenStudy (anonymous):

Thank you!

OpenStudy (mayankdevnani):

@thivitaa

OpenStudy (anonymous):

4x=-8 x=-2

OpenStudy (anonymous):

thanks.